Mr. R has bought call and put options. The contract consists of 100 shares. Mr. R purchased a 3-month call option with a strike price of $57 and $2 as premium. He paid $2 per share as premium for buying 3-month put option with a strike price of $55. What will be the situation of Mr. R if the stock price increases to $59 in 3 month or if stock price reduces to $50 in 3 months
1
Expert's answer
2020-06-21T17:59:20-0400

If stock price increases to $59, Mr. R get nothing from call option "(57+2-59)*100=0"

If stock price increases to $59, Mr. R get 200$ from put option "(59-55-2)*100=200"

If stock price reduces to $50, Mr. R get 900$ from call option "(57+2-50)*100=900"

If stock price reduces to $50, Mr. R loss 700$ from put option "(50-55-2)*100=-700"